OpenStudy (anonymous):

A distribution center receives shipments from three different factories in quantities of 50, 35, and 20. Three times a product is selected at random. Fing the probability that none of the products came from the second factory

OpenStudy (kropot72):

Assuming the sampling is done without replacement: On the first selection P(0) = 70/105 On the second selection P(0) = 69/104 On the third selection P(0) = 68/103 \[P(0\ from\ 3\ selections)=\frac{70\times 69\times 68}{105\times 104\times 103}=you\ can\ calculate\]